OpenStudy (anonymous):

Which function transforms the graph of y=x^2 so that it is first shifted up 3 units and is then reflected across the x-axis? A. y=(-x)^2+3 B. y=-(x+3)^2 C. y=-x^2-3 D y=-x^2+3

OpenStudy (amistre64):

since it aint shifted left or right; there is no (x-...)^2 to worry about

OpenStudy (amistre64):

i assume A is a typo

OpenStudy (anonymous):

no its right

OpenStudy (amistre64):

if we move it up 3 then flip it over the x axis; we go to -3 my gut says C

OpenStudy (anonymous):

To 'reflect across x' you multiply the y part by -1

OpenStudy (anonymous):

i feel like its B, but not positive

OpenStudy (anonymous):

Starting from the beginning we have: y = x^2 We shift it up 3: y - 3 = x^2 We reflect across x: -(y - 3) = x^2 Now that is one of your answers, though their version is a different form of the same thing because they moved some of the bits around to solve for y.

OpenStudy (curry):

amis

OpenStudy (curry):

help needed